OpenStudy (anonymous):

I have to solve this using linear equations. But I don't understand how to model the equation with the information given. Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April? A. 27 B. 35 C. 95 D. 103

OpenStudy (neer2890):

Let Jeannie's age is x years and April's age is y years. so Jeannie is 8 years older than her younger sister April i.e.,x=y+8 When added together, their ages equal 198 i.e., x+y=198

OpenStudy (neer2890):

solve the two equations and find their ages.

OpenStudy (anonymous):

April is 95 years old? is that right?

OpenStudy (neer2890):



OpenStudy (neer2890):

correct..!

OpenStudy (anonymous):

I used substitution to solve, not elimination. Either works. But we already had x=y+8 so I just plugged that into x+y=198.

OpenStudy (neer2890):

in mathematics, all that matters is to find correct results eitherway.

OpenStudy (anonymous):

Can you help me with another? I still don't understand how to model and equation with the age differences.

OpenStudy (neer2890):

ok

OpenStudy (anonymous):

Rachel is 3 years older than Sarah. Their ages add to 67. I have to find the correct system..

OpenStudy (neer2890):

Let the ages of Rachel and Sarah are x and y years respectively. Rachel is 3 years older than Sarah means Age of rachel(x)= Age of Sarah(y)+3 so x=y+3 Their ages add to 67 i.e., x+y=67

OpenStudy (anonymous):

Oh thank you! I get it now.

OpenStudy (neer2890):

you're welcome..:)